# A Survey on Text-Driven 360-Degree Panorama Generation

### v2

@@ Abstract
The advent of text-driven 360-degree panorama generation, enabling the synthesis of 360-degree panoramic images directly from textual descriptions, marks a transformative advancement in immersive visual content creation. This innovation significantly simplifies the traditionally complex process of producing such content. Recent progress in text-to-image diffusion models has accelerated the rapid development in this emerging field. This survey presents a comprehensive review of text-driven 360-degree panorama generation, offering an in-depth analysis of state-of-the-art algorithms and their expanding applications in 360-degree 3D scene generation. Furthermore, we critically examine current limitations and propose promising directions for future research. A curated project page with relevant resources and research papers is available at https://littlewhitesea.github.io/Text-Driven-Pano-Gen/.
Index Terms:
360-degree panorama generation, text-driven generation, text-to-image diffusion model, 360-degree 3D scene generation.I Introduction
Rapid growth of immersive technologies, such as virtual reality (VR) and augmented reality (AR), has dramatically increased the demand for high-quality panoramic visual content. Among such content, 360-degree panoramas are pivotal in delivering realistic and immersive experiences by capturing a complete spherical view of an environment. Traditionally, producing these panoramas requires specialized camera equipment and considerable technical expertise. However, recent advances in text-driven 360-degree panorama generation [14, 40, 56, 57] have introduced groundbreaking capabilities, enabling the synthesis of 360-degree panoramic images directly from textual descriptions. This innovation can potentially revolutionize content creation over diverse domains including VR/AR applications, gaming, and virtual tours.
@@
The availability of large-scale paired image-text datasets has facilitated the development of text-to-image latent diffusion models (LDMs) [63], which excel at synthesizing high-quality, visually compelling images aligned with given text descriptions [64, 65, 66, 67, 72]. Leveraging the powerful generative capabilities of pre-trained LDMs, researchers have developed methods specifically tailored to address the unique challenges of text-driven 360-degree panorama generation [42, 57, 58, 59, 60, 73]. Despite the rapid progress in this field, there is currently no survey that provides a systematic analysis of these emerging methodologies, elucidates their key characteristics, and highlights their applications. To address this gap, this paper presents a holistic survey and analysis of the latest research advances in text-driven 360-degree panorama generation.
This survey is structured as follows: First, we establish a foundational understanding of this field by introducing the principal representations of 360-degree panoramas, presenting prominent datasets commonly used in this area, and outlining key evaluation metrics employed to assess the quality and fidelity of generated panoramic content. Next, we review state-of-the-art (SOTA) methods for text-driven 360-degree panorama generation, categorizing them into two primary paradigms: (a) Text-Only Generation and (b) Text-Driven Narrow Field-of-View (NFoV) Outpainting. Fig. 2 and Fig. 3 provide a systematic taxonomy and a chronological overview of these SOTA methods, respectively. Methods on text-only generation aim to synthesize entire 360-degree panoramas exclusively from textual prompts, typically utilizing fine-tuning techniques such as Low-Rank Adaptation (LoRA) [39] or DreamBooth [38] on pre-trained LDMs [63]. In contrast, text-driven NFoV outpainting allows greater user control by conditioning the generation process on both textual descriptions and initial NFoV images. Following this, we explore emerging applications of text-driven 360-degree panorama generation in the realm of 360-degree 3D scene generation. Finally, we discuss the prevailing challenges in this developing field and propose potential directions for future research.
This paper offers the first survey on text-driven 360-degree panorama generation, systematically reviewing its state-of-the-art techniques, key datasets and evaluation metrics, as well as exploring its applications in 360-degree 3D scene generation. Furthermore, we identify critical challenges and outline future research directions, aiming to offer a valuable resource to researchers and practitioners in this area.
@@ II-A Text-to-Image Diffusion Models
Recently, 3D Gaussian Splatting (3DGS) [26, 35] has emerged as an efficient alternative to 3D scene representation. By combining an explicit representation of 3D Gaussians with a highly efficient differentiable rasterization pipeline, 3DGS facilitates rapid scene reconstruction and rendering. This advancement has opened up new possibilities, including recent explorations in text-to-3D 360-degree scene synthesis [33, 32], which leverage text-driven 360-degree panorama generation techniques.
@@ III-A Representations of 360-Degree Panoramas
The representation of 360-degree panoramic content poses a fundamental challenge: How to accurately map spherical visual information onto a two-dimensional plane? To address this, a variety of projection methodologies [7] have been developed, each with distinct advantages and trade-offs. Below, we first outline two widely used formats for 360-degree panorama representation: Equirectangular Projection (ERP) and Cubemap Projection (CMP), as illustrated in Fig. 4, and then introduce Multi-Perspective Projection (MPP), a category we define for the classification purposes of this survey.
@@ III-B Datasets
360-degree panoramic image generation from text prompts presents unique challenges due to the complete 360180 field of view that these images encompass. Text-to-image diffusion models [12, 21, 63], predominantly trained on perspective images with a narrower field of view, often struggle to synthesize high-quality 360-degree panoramas. To address this, several specialized datasets have been developed to facilitate research in this domain. Tab. I summarizes these datasets, with further details provided below.
@@ III-C Evaluation Metrics
IS [54] measures both the quality and diversity of generated images by leveraging Inception-v3 [36]. It calculates the KL divergence between the conditional class distribution of generated images and the marginal distribution over all generated samples. Higher IS scores suggest better visual quality and diversity.
CS [52] evaluates consistency between text prompts and generated images using the CLIP model [52]. It calculates the cosine similarity between the text embedding of the prompt and the visual embedding of the generated image. A higher CS score reflects stronger text-image alignment and semantic coherence.
@@
UniPano [80] | arxiv 2025 | Matterport3D [49] | ERP |
The universal metrics provide general assessments but often fail to capture nuances and geometric distortions unique to 360-degree panoramas. Panorama-specific metrics designed specifically for ERP-based 360-degree panoramic images include Fréchet Auto-Encoder Distance (FAED), Omnidirectional FID (OmniFID), Discontinuity Score (DS), and Distortion-perception FID (Distort-FID).
@@
DS [51] measures the seam alignment across the borders of generated panoramas by applying a kernel-based edge detection algorithm. Lower discontinuity scores correspond to fewer visible seam artifacts, indicating better perceived consistencies across the seam.
@@
We provide a comprehensive comparison of state-of-the-art methods, introduced in the following section, using the outlined metrics in Sec. IV-C2.
IV State-of-the-Art Methods
@@ IV-A1 Training-Based
Text2Light [56], an early notable effort, explored a hierarchical text-driven framework utilizing VQGAN [43] and CLIP [52] to address this challenge based on training data aggregated from multiple sources, such as Polyhaven [5], Laval Indoor [45] and Laval Outdoor [46]. Recently, the advent of latent diffusion models (LDMs) [63] for text-to-image synthesis marked a significant advancement, enabling more sophisticated 360-degree panorama generation techniques. LDMs are typically trained on vast datasets consisting of standard perspective images with a limited field of view and corresponding textual descriptions. Despite demonstrating robust capabilities in generating perspective images from text prompts, these models face significant difficulties when creating 360-degree panoramas with a complete 360180 field of view, which differ substantially from traditional perspective images.
@@
To adapt pre-trained LDMs for 360-degree panorama synthesis, a common strategy is to fine-tune these models with specialized 360-degree panorama datasets. Diffusion360 [58] exemplifies this approach by leveraging the DreamBooth technique [38] to fine-tune a pre-trained LDM [63] on SUN360 [48]. To ensure geometric consistency of boundaries, Diffusion360 utilizes a circular blending strategy during both the denoising process and the VAE [74] decoding stage, effectively reducing seam artifacts. In addition, it introduces a super-resolution module to enable the generation of high-resolution (61143072) 360-degree panoramas. Notably, the DreamBooth fine-tuning approach is computationally demanding and time-consuming because it generally fine-tunes a significant portion of the model parameters.
@@
In contrast, LoRA [39] has recently gained attention as a parameter-efficient fine-tuning method. LoRA works by injecting trainable low-rank matrices into the pre-trained model’s weights, allowing for rapid adaptation to new tasks with minimal additional parameters. For example, StitchDiffusion [57] employs LoRA to fine-tune a pre-trained LDM on a dataset of paired 360-degree images (sourced from Polyhaven [5]) and corresponding textual descriptions generated using BLIP [17]. It further introduces a tailored generation method based on MultiDiffusion [55] to ensure boundary continuity in the generated panoramas.
@@
[23] | arxiv 2025 | SUN360 [48], Structured3D [47] | MPP | |||
PAR [25] | arxiv 2025 | Matterport3D [49] | ERP |
Other works have similarly adopted the LoRA fine-tuning technique. PanFusion [59] devises a dual-branch diffusion model, trained on the Matterport3D [49] dataset, with separate LoRA layers to integrate both global panoramic and local perspective views. It introduces an equirectangular-perspective projection attention module to facilitate information exchange between the two branches, aiming to alleviate visual inconsistencies in the generated 360-degree panoramas. However, PanFusion’s output often exhibits blurriness at the top and bottom regions, due to its training dataset. To avoid this issue, DiffPano [40] utilizes the Habitat Matterport 3D dataset [15] to produce multi-view consistent 360-degree panoramas with clearer top and bottom details. For generating more precise textual descriptions, DiffPano adopts BLIP2 [18] and Llama2 [16] sequentially, resulting in a panoramic video-text dataset. Based on this dataset, it fine-tunes a pre-trained LDM [63] using LoRA for single-view text-driven 360-degree panorama generation. Furthermore, to enable multi-view 360-degree panorama generation based on text prompts and camera viewpoints, DiffPano proposes a spherical epipolar-aware attention module, which effectively extends single-view panorama generation to multi-view consistent panorama generation. Based on the observation that value and output weight matrices are crucial during the LoRA-based fine-tuning of cross-attention blocks within pre-trained LDMs for 360-degree panoramic image generation, UniPano [80] proposes a uni-branch framework for panorama generation. This framework exclusively fine-tunes these specific matrices using LoRA, while keeping the original query and key weight matrices in the cross-attention blocks frozen.
More recently, diverging from fine-tuning the commonly used pre-trained LDMs [63] (as in PanFusion [59] and DiffPano [40]), several studies have explored alternative pre-trained models for synthesizing 360-degree panoramic images from textual descriptions. Inspired by OmniGen [76], [23] adopted a diffusion model consisting of a VAE [74] and a pre-trained Transformer [75]. To adapt the pre-trained Transformer for synthesizing 360-degree panoramic images, LoRA fine-tuning technique is employed. Instead of processing the entire 360-degree panoramic image at once, generates six overlapping viewports. In the inference phase, these synthesized perspective images are integrated to reconstruct 360-degree panoramic images. Addressing the challenge that spatial distortions in ERP 360-degree panoramic images violate the identically and independently distributed (i.i.d.) Gaussian noise assumption inherent in many diffusion models, PAR [25], inspired by masked autoregressive modeling (MAR) [77], proposes an autoregressive modeling approach for text-based 360-degree panoramic image generation. This method is not constrained by the i.i.d. assumption. Specifically, PAR fine-tunes a pre-trained autoregressive model [78] on the Matterport3D dataset [49] and develops a dual-space circular padding technique to mitigate boundary discontinuities.
IV-A2 Training-Free
In contrast to training-based approaches, PanoFree [41], introduces a tuning-free multi-view image generation framework based on a pre-trained LDM [63]. Guided by textual descriptions, PanoFree leverages iterative warping and inpainting steps to produce multi-view perspective images, which are subsequently stitched into 360-degree panoramas, thus avoiding the need for specialized 360-degree panorama datasets. Unlike PanoFree, which operates on perspective latent representations, the recent SphereDiff [22] proposes leveraging a spherical latent representation for training-free generation of 360-degree panoramic images. Specifically, SphereDiff extends the MultiDiffusion [55] framework to this spherical latent space. To mitigate minor distortions arising from the spherical-to-perspective projection during its process, SphereDiff further incorporates a distortion-aware weighted averaging method.

### v3

@@ Abstract
The advent of text-driven 360-degree panorama generation, enabling the synthesis of 360-degree panoramic images directly from textual descriptions, marks a transformative advancement in immersive visual content creation. This innovation significantly simplifies the traditionally complex process of producing such content. Recent progress in text-to-image diffusion models has accelerated the rapid development in this emerging field. This survey presents a comprehensive review of text-driven 360-degree panorama generation, offering an in-depth analysis of state-of-the-art algorithms. We extend our analysis to two closely related domains: text-driven 360-degree 3D scene generation and text-driven 360-degree panoramic video generation. Furthermore, we critically examine current limitations and propose promising directions for future research. A curated project page with relevant resources and research papers is available at https://littlewhitesea.github.io/Text-Driven-Pano-Gen/.
I Introduction
Rapid growth of immersive technologies, such as virtual reality (VR) and augmented reality (AR), has dramatically increased the demand for high-quality panoramic visual content. Among such content, 360-degree panoramas are pivotal in delivering realistic and immersive experiences by capturing a complete spherical view of an environment. Traditionally, producing these panoramas requires specialized camera equipment and considerable technical expertise. However, recent advances in text-driven 360-degree panorama generation [14, 40, 56, 57, 58, 59] have introduced groundbreaking capabilities, enabling the synthesis of 360-degree panoramic images directly from textual descriptions. This innovation not only revolutionizes content creation over diverse domains [85, 86, 87, 89] including VR/AR applications, gaming, and virtual tours, but also serves as a foundational technology for new creative frontiers.
@@ I Introduction
The availability of large-scale paired image-text datasets has facilitated the development of text-to-image latent diffusion models (LDMs) [63], which excel at synthesizing high-quality, visually compelling images aligned with given text descriptions [64, 65, 66, 67, 72]. Leveraging the powerful generative capabilities of pre-trained LDMs, researchers have developed methods specifically tailored to address the unique challenges of text-driven 360-degree panoramic image generation [42, 57, 58, 59, 60, 73]. Although broader surveys on panoramic vision and 3D scene-generation [88, 104] briefly discuss some text-driven 360-degree panorama generation methods, they treat them only as peripheral topics.
To the best of our knowledge, a focused and systematic analysis devoted specifically to text-driven 360-degree panoramic image generation has not yet been presented. To address this gap, this paper presents a holistic survey and analysis of text-driven 360-degree panorama generation, its direct applications, and related emerging fields.
This survey is structured as follows: First, we establish a foundational understanding of this field by introducing the principal representations of 360-degree panoramas, presenting prominent datasets commonly used in this area, and outlining key evaluation metrics employed to assess the quality and fidelity of generated panoramic content. Next, we review state-of-the-art (SOTA) methods for text-driven 360-degree panorama generation, categorizing them into two primary paradigms: (a) Text-Only Generation and (b) Text-Driven Narrow Field-of-View (NFoV) Outpainting. Fig. 2 and Fig. 3 provide a systematic taxonomy and a chronological overview of these SOTA methods, respectively. Following this, we explore two key emerging directions that are closely related to this field: (a) text-driven 360-degree 3D scene generation, which uses 360-degree panoramic images as an intermediate step; and (b) text-driven 360-degree panoramic video generation, a parallel and more complex task. Finally, we discuss the prevailing challenges in this developing field and propose potential directions for future research.
In short, this paper offers the first dedicated and comprehensive survey on text-driven 360-degree panoramic image generation, systematically reviewing its state-of-the-art techniques, key datasets and evaluation metrics. Furthermore, we explore its two closely related emerging directions: text-driven 360-degree 3D scene generation and text-driven 360-degree panoramic video synthesis. We also identify critical challenges and outline future research directions, aiming to offer a valuable resource to researchers and practitioners in this area.
@@ II-A Text-to-Image Diffusion Models
Recently, 3D Gaussian Splatting (3DGS) [26] has emerged as an efficient alternative to 3D scene representation. By combining an explicit representation of 3D Gaussians with a highly efficient differentiable rasterization pipeline, 3DGS facilitates rapid scene reconstruction and rendering. This advancement has opened up new possibilities, including recent explorations in text-to-3D 360-degree scene synthesis [32, 33], which leverage text-driven 360-degree panorama generation techniques.
@@ III-A Representations of 360-Degree Panoramas
The representation of 360-degree panoramic content poses a fundamental challenge: How to accurately map spherical visual information onto a two-dimensional plane? To address this, a variety of projection methodologies [7, 88] have been developed, each with distinct advantages and trade-offs. Below, we first outline two widely used formats for 360-degree panorama representation: Equirectangular Projection (ERP) and Cubemap Projection (CMP), as illustrated in Fig. 4, and then introduce Multi-Perspective Projection (MPP), a category we define for the classification purposes of this survey.
@@ III-A3 Multi-Perspective Projection (MPP)
The advantages and limitations of these representation formats are further analyzed in Sec. IV-C1.
@@ III-B Datasets
360-degree panoramic image generation from text prompts presents unique challenges due to the complete 360180 field of view that these images encompass. Text-to-image diffusion models [12, 21, 35, 63], predominantly trained on perspective images with a narrower field of view, often struggle to synthesize high-quality 360-degree panoramas. To address this, several specialized datasets have been developed to facilitate research in this domain. Tab. I summarizes these datasets, with further details provided below.
@@ III-C Evaluation Metrics
IS [54] measures both the quality and diversity of generated images by leveraging Inception-v3 [36]. It calculates the KL divergence between the conditional class distribution of generated images and the marginal distribution over all generated samples. Higher IS suggests better visual quality and diversity.
CS [52] evaluates consistency between text prompts and generated images using the CLIP model [52]. It calculates the cosine similarity between the text embedding of the prompt and the visual embedding of the generated image. A higher CS reflects stronger text-image alignment and semantic coherence.
@@
| UniPano [80] | ICCV 2025 | Matterport3D [49] | ERP | |||
| SMGD [81] | CVPR 2025 | Matterport3D [49] | Spherical |
However, these universal metrics have significant limitations when applied to 360-degree panoramic content. The core issue lies in their underlying encoders (e.g., Inception-v3, CLIP’s ViT), which were trained primarily on large datasets of standard perspective images. These encoders are not trained to account for the geometric distortions inherent in equirectangular projections. Consequently, FID, KID, and IS may penalize geometrically correct panoramic features as artifacts, leading to an inaccurate assessment of perceptual quality. Similarly, the CS may fail to accurately measure text-image alignment, as its encoder might misinterpret distorted objects or spatial relationships within the spherical scene. This gap highlights the inadequacy of universal metrics for capturing properties unique to 360-degree panoramas, such as seamlessness and global geometric fidelity. To address these shortcomings, several panorama-specific metrics have been proposed, including Fréchet Auto-Encoder Distance (FAED), Omnidirectional FID (OmniFID), Discontinuity Score (DS), and Distortion-perception FID (Distort-FID).
@@
DS [51] measures the seam alignment across the borders of generated panoramas by applying a kernel-based edge detection algorithm. A lower DS corresponds to fewer visible seam artifacts, indicating better perceived consistencies across the seam.
@@
Despite these advances, a critical gap remains. While metrics like OmniFID improve the measurement of geometric fidelity by mitigating ERP distortions and DS checks for visual seamlessness, they do not holistically evaluate the global semantic and structural coherence of a full 360-degree panoramic scene. Developing new metrics that capture this global consistency is therefore an important direction for future research.
We provide a comprehensive comparison of state-of-the-art methods, introduced in the following section, using the outlined metrics in Sec. IV-D2.
IV State-of-the-Art Methods for Images
@@ IV-A1 Training-Based
Text2Light [56], an early notable effort, explores a hierarchical text-driven framework, using VQGAN [43] and CLIP [52] to address this challenge based on training data aggregated from multiple sources, such as Polyhaven [5], Laval Indoor [45] and Laval Outdoor [46]. Recently, the advent of latent diffusion models (LDMs) [63] for text-to-image synthesis marks a significant advancement, enabling more sophisticated 360-degree panorama generation techniques. LDMs are typically trained on vast datasets consisting of standard perspective images with a limited field of view and corresponding textual descriptions. Despite demonstrating robust capabilities in generating perspective images from text prompts, these models face significant difficulties when creating 360-degree panoramas with a complete 360180 field of view, which differ substantially from traditional perspective images.
@@
To adapt pre-trained LDMs for 360-degree panorama synthesis, a common strategy is to fine-tune these models with specialized 360-degree panorama datasets. Diffusion360 [58] exemplifies this approach by leveraging the DreamBooth technique [38] to fine-tune a pre-trained LDM [63] on SUN360 [48]. To ensure geometric consistency of boundaries, Diffusion360 uses a circular blending strategy during both the denoising process and the VAE [74] decoding stage, effectively reducing seam artifacts. In addition, it introduces a super-resolution module to enable the generation of high-resolution (61143072) 360-degree panoramas. While this full fine-tuning approach adopted by Diffusion360 effectively embeds panoramic geometry into the model, its primary trade-off is the high computational cost and risk of quality drop by altering the model’s original generative priors.
@@
In contrast, LoRA [39] has recently gained attention as a parameter-efficient fine-tuning method. LoRA works by injecting trainable low-rank matrices into the pre-trained model’s weights, allowing for rapid adaptation to new tasks with minimal additional parameters. For example, StitchDiffusion [57] employs LoRA to fine-tune a pre-trained LDM on a dataset of 120 paired 360-degree images (sourced from Polyhaven [5]) and corresponding textual descriptions generated using BLIP [17]. Its key contribution is formulating panorama generation as a latent-space stitching problem, using a MultiDiffusion-based [55] method to enforce boundary continuity. However, the geometry fidelity of 360-degree panoramas generated by StitchDiffusion is relatively low due to the small fine-tuning dataset.
Other works [40, 59, 80, 84] have similarly adopted the LoRA fine-tuning technique. PanFusion [59] contributes a novel dual-branch architecture, trained on the Matterport3D [49] dataset, with separate LoRA layers to integrate both global panoramic and local perspective views. It introduces an equirectangular-perspective projection attention module to facilitate information exchange between the two branches, aiming to alleviate visual inconsistencies in the generated 360-degree panoramas. However, PanFusion’s output often exhibits blurriness at the top and bottom regions, due to its training dataset. To avoid this issue, DiffPano [40] uses the Habitat Matterport 3D dataset [15] to produce multi-view consistent 360-degree panoramas with clearer top and bottom details. For generating more precise textual descriptions, DiffPano adopts BLIP2 [18] and Llama2 [16] sequentially, resulting in a panoramic video-text dataset. Based on this dataset, it fine-tunes a pre-trained LDM [63] using LoRA for single-view text-driven 360-degree panorama generation. Furthermore, to enable multi-view 360-degree panorama generation based on text prompts and camera viewpoints, DiffPano introduces a spherical epipolar-aware attention module, a key innovation for enforcing multi-view geometric consistency. Based on the observation that value and output weight matrices are crucial during the LoRA-based fine-tuning of cross-attention blocks within pre-trained LDMs for 360-degree panoramic image generation, UniPano [80] proposes a uni-branch framework for panorama generation. This framework exclusively fine-tunes these specific matrices using LoRA, while keeping the original query and key weight matrices in the cross-attention blocks frozen.
Diverging from fine-tuning the commonly used pre-trained LDMs [63] (as in PanFusion [59] and DiffPano [40]), several studies have explored alternative pre-trained models for synthesizing 360-degree panoramic images from textual descriptions. Inspired by OmniGen [76], [23] adopts a diffusion model consisting of a VAE [74] and a pre-trained Transformer [75]. To adapt the pre-trained Transformer for synthesizing 360-degree panoramic images, the LoRA fine-tuning technique is employed. Instead of processing the entire 360-degree panoramic image at once, generates six overlapping viewports. In the inference phase, these synthesized perspective images are integrated to reconstruct 360-degree panoramic images. Addressing the challenge that spatial distortions in ERP 360-degree panoramic images violate the identically and independently distributed (i.i.d.) Gaussian noise assumption inherent in many diffusion models, PAR [25], inspired by masked autoregressive modeling (MAR) [77], proposes an autoregressive modeling approach for text-based 360-degree panoramic image generation. This method is not constrained by the i.i.d. assumption. Specifically, PAR fine-tunes a pre-trained autoregressive model [78] on the Matterport3D dataset [49] and develops a dual-space circular padding technique to mitigate boundary discontinuities.
Most aforementioned approaches rely on ERP representations, which struggle to adequately deal with the inherent spherical distortions. To mitigate these distortions and maintain global geometric coherence, SMGD [81] proposes the use of spherical manifold convolution within a spherical manifold U-Net combined with VQGAN [43], enabling more accurate synthesis of 360-degree panoramic images, but the primary trade-off is reduced transferability of pre-trained diffusion priors, since the specialized spherical convolutions are not directly compatible with standard text-to-image diffusion architectures [63, 21].
IV-A2 Training-Free
In contrast to training-based approaches, training-free methods avoid any model fine-tuning and instead repurpose powerful pre-trained text-to-image diffusion backbones. PanoFree [41] pioneered a tuning-free multi-view image generation framework based on a pre-trained LDM [63]. Guided by textual descriptions, PanoFree leverages iterative warping and inpainting steps to produce multi-view perspective images, which are subsequently stitched into 360-degree panoramas, thus avoiding the need for specialized 360-degree panorama datasets. While this avoids the need for specialized training data, its multi-step process can be slow and risks accumulating errors that harm global consistency. Unlike PanoFree, which operates on perspective latent representations, the recent SphereDiff [22] contributes a more theoretically grounded approach by extending the MultiDiffusion [55] framework to a spherical latent space. To mitigate minor distortions arising from the spherical-to-perspective projection during its process, SphereDiff further incorporates a distortion-aware weighted averaging method. Although these approaches inherit the rich prior knowledge of large text-to-image models and require no additional training or 360-degree panoramic data, their patch-based synthesis mechanisms can lead to global inconsistencies and comparatively longer inference times than training-based models. Future work may explore stronger global guidance and more efficient inference designs to overcome these limitations.
@@ IV-A2 Training-Free
| [23] | ACM MM 2025 | SUN360 [48], Structured3D [47] | MPP | |||
| PAR [25] | arxiv 2025 | Matterport3D [49] | ERP | |||
| SpND [82] | ICML 2025 | Matterport3D [49], Structured3D [47] | ERP | |||
| DreamCube [83] | ICCV 2025 | Structured3D [47] | CMP |